# Vevo2: A Unified and Controllable Framework for Speech and Singing Voice Generation

### v2

@@ I Introduction
Controllable generation of human voice is an important research topic in the field of audio generation [tts-book-tanxu, xietianxin-survey]. In recent years, remarkable progress has been achieved in speech generation, particularly in zero-shot text-to-speech (TTS) [naturalspeech3, seedtts, maskgct, cosyvoice], largely propelled by the availability of large-scale training corpora. However, for more challenging domains such as singing voice—which can be regarded as a highly expressive form of speech with strictly constrained prosody (e.g., melody) [^1] This paper adopts a definition in which prosody refers to the suprasegmental attributes of human voice, including pitch, duration, and loudness. We regard melody—which primarily involves pitch and duration (with rhythm viewed as a temporal organization of duration)—as a specific and more constrained form of prosody. [singingvoice-science]—achieving both high generation quality and controllability remains challenging [seedmusic, yue].
@@
This paper focuses on unifying controllable speech and singing voice generation. Our motivation is based on the hypothesis that learning for speech and singing voice can be mutually beneficial through unified modeling. Specifically, (1) Speech $\rightarrow$ Singing Voice: the abundance of speech data can supplement the typically smaller singing voice datasets [emilia, gtsinger], potentially improving the overall quality of generated singing voice [yue]. (2) Singing Voice $\rightarrow$ Speech: the inherent expressiveness of singing voice, when incorporated into a unified learning framework, can enhance expressive speech generation [yicheng-trans]. Besides, since singing voices inherently contain structured melodic patterns, learning to model them helps the system enhance its ability to follow melody, thereby improving overall prosody-following capability. This capability is particularly beneficial for tasks such as prosody-preserved speech editing (Section V-B).
@@
Several recent studies have attempted to unify speech and singing voice generation. For instance, UniSyn [unisyn] integrates TTS and Singing Voice Synthesis (SVS) through a task identifier that switches between generation modes, while UniAudio [uniaudio] extends this concept to a broader range of audio generation tasks via task-conditioned sequence modeling. Although these approaches represent valuable progress, they still face two fundamental challenges. First, the existing corpora for singing voice generation rely heavily on expert annotations [singing-expression, diffsinger, tcsinger, spsinger], such as detailed music notation with precise alignment between lyrics and musical notes [m4singer, singstyle111, gtsinger]. These datasets are scarce, and their unique data organization is inherently inconsistent with large-scale speech corpora, thereby limiting scalability and unified training. Second, achieving versatile controllability within such a unified system remains nontrivial—such as the independent manipulation of text (or lyrics), prosody (or melody), style (e.g., accent, emotion, or singing techniques [vocalset, gtsinger]), and timbre (i.e., speaker identity) within a single framework.
@@
Voice Conversion (style-converted)
source wav, reference wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Singing Voice Conversion (style-converted)
$\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
@@
Voice Conversion (style-preserved)
source wav, reference wav - $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$]
@@
To address these challenges, this paper introduces Vevo2, a versatile and controllable voice generation framework designed to unify speech and singing voice. Vevo2 adopts a two-stage architecture, similar to classic zero-shot speech generation systems [seedtts, cosyvoice, vevo] (Figure 1): (1) Auto-Regressive (AR) Content-Style Modeling Stage: given a target text and a prosodic source as input, we utilize an AR transformer to generate the content-style tokens, which is prompted by a style reference. (2) Flow-Matching (FM) Acoustic Modeling Stage: given content-style tokens as input, we employ an FM transformer to produce acoustic representations (specifically, Mel spectrograms in our study), which is prompted by a timbre reference. To achieve the desired unification and controllability, Vevo2 introduces the following three key method designs:
1. Unified Speech and Singing Voice Tokenizers. We propose two distinct audio tokenizers: a unified prosody tokenizer and a unified content-style tokenizer. (1) Prosody Tokenizer: A single-codebook VQ-VAE tokenizer [vq-vae, rvqgan] operating at 6.25 Hz (56.25 bps), trained to reconstruct the chromagram [meinard-book] of raw audio (Figure 2(a)). This tokenizer unifies the modeling of both melody in singing voices and prosody in speech, and can directly extract tokens from the chromagram of raw audio, eliminating the need for expert-annotated music notation and thus greatly improving scalability. Moreover, due to the chromagram’s strong generalization ability across various types of audio data [meinard-book, musicgen], this tokenizer can even capture the melody of non-human sounds such as instrumental music (Section V-F), thereby enabling unique music generation applications, such as humming-to-singing and instrument-to-singing. (2) Content-Style Tokenizer: It employs a similar architecture to the prosody tokenizer, but operates at 12.5 Hz (175 bps) and is trained to reconstruct both chromagram features and Whisper hidden features [whisper] (Figure 2(b)). It effectively captures linguistic content, melody, and style in both speech and singing voice, while also achieving robust timbre disentanglement (Table III).
@@
3. Multi-objective Alignment. To further improve the text- and prosody-following abilities of the pre-trained model, we introduce a multi-objective post-training task. We propose a composite reward that incorporates both intelligibility and prosody similarity, and adopt the Group Relative Policy Optimization (GRPO) algorithm [grpo] for training. Experimental results demonstrate that this post-training significantly enhances the model’s controllability with respect to text and melody, and also improves its generalization to out-of-pretraining-distribution data, such as instrumental sounds.
@@
We will release code and model checkpoints of this study at Amphion [amphion]. [^2] A preliminary version of this work was released as a preprint on arXiv [vevo2-arxiv].
@@ II Related Work
Controllable Voice Generation Controllable speech and singing voice generation has consistently remained a significant research focus. (1) In the speech domain, existing works explore control over multiple attributes: linguistic content, which is fundamental to TTS [tts-book-tanxu, xietianxin-survey] and speech editing tasks [voicebox, voicecraft, ssrspeech, f5tts]; timbre, which is central to VC tasks [ppg-vc, autovc, naturalspeech3, voiceshop, vevo, seedvc]; and style, which encompasses accent [parallel-ac-zhaoguanlong21, convertandspeak, vevo], emotion [esd, pavits, vevo], whisper [whisper-to-normal-lstm], and a broader natural language description [parler-tts, controlspeech, paraspeechcaps]. (2) For singing voice, beyond controlling lyrics, timbre, and style—as explored in SVS [diffsinger, visinger, tcsinger, stylesinger, techsinger], SVC [svcc-2023, svcc-vits-ziqian, neucosvc, amphion-svc], and singing lyric editing tasks [editsinger, unisinger, songeditor]melody control holds particular significance due to the inherent musicality of singing voices [musicgen, seedmusic, yue]. (3) From a unified modeling perspective, previous research has attempted to integrate various speech generation tasks such as TTS and VC [vevo, cosyvoice, metis], or different singing voice generation tasks like SVS and SVC [unisinger, everyone-can-sing]. However, models capable of unifying both speech and singing voice domains remain scarce [uniaudio, unisyn]. Moreover, the potential benefits of joint speech and singing pre-training, along with the versatile control mechanisms for multiple voice attributes within a unified system, remain largely unexplored.
@@
Speech Tokenizer Recent research has emphasized the disentangled speech representation learning, aiming to separately model the factorized speech attributes to reduce learning complexity [megatts, naturalspeech3] and achieve more precise controllability [speech-resynthesis-interspeech21, naturalspeech3, vevo]. Common speech tokenizers include linguistic content tokenizers [hubert, vevo, syllablelm], content-style tokenizers [audiolm, seedtts, maskgct, cosyvoice, vevo], and prosody tokenizers [megatts, naturalspeech3]. However, in pursuing a unified controllable model for both speech and singing generation, existing speech tokenizers exhibit several limitations: As these tokenizers predominantly focus on the speech domain, their effectiveness in handling singing voice data remains unexplored. The design of a unified prosody tokenizer capable of handling both modalities, along with enhancing the melody encoding capabilities of content-style tokenizers [amphion-svc], remains an important issue for future exploration.
@@
Post-training for Audio Generation Alignment via post-training has demonstrated its effectiveness in the generation of text [instructgpt, RLHF-anthoropic], vision [imagereward, tldr-perturb-vision], speech [speechalign, seedtts, cosyvoice2], music [musicrl], and sound effects [tango2, baton]. In speech generation, existing works have employed preference alignment to enhance multiple aspects of speech, including intelligibility [seedtts, cosyvoice2, intp], speaker similarity [seedtts, cosyvoice2], emotion controllability [seedtts, emo-dpo], and overall quality [speechalign, uno, fpo]. For developing a unified speech and singing voice generation model, the importance of specific capability elicitation through post-training becomes particularly crucial, given the expanded range of capabilities injected into the pre-trained model. However, the design of appropriate alignment objectives for unified speech and singing voice generation remains an open research problem.
@@ III Methodology
The overall architecture of Vevo2 is a two-stage, versatile voice generation framework that unifies both speech and singing voice (Figure 1). During inference, by flexibly combining different input conditions—such as text, prosodic source, style reference, and timbre reference—Vevo2 supports a wide range of voice-related synthesis, conversion, and editing tasks (Section III-D). Specifically, Vevo2 first encodes linguistic, stylistic, and prosodic cues into discrete content-style tokens through an AR transformer, and then reconstructs Mel spectrograms via an FM transformer. These spectrograms are subsequently converted into waveforms using a Vocos-based vocoder [vocos]. To achieve the unified speech-singing voice generation and versatile controllability, Vevo2 is built upon several key methodological designs, which are detailed in the following sections.
@@ III-A Unified Speech and Singing Voice Tokenizers
Prosody Tokenizer In order to model both the prosody of speech and the melody of singing voice using a unified method, we choose chromagram feature [meinard-book] and design a VQ-VAE tokenizer [vq-vae, rvqgan] based on it. Some previous studies of speech generation tried to use fundamental frequency (F0) [speech-resynthesis-interspeech21, naturalspeech3] or the low-frequency band of the mel spectrogram [megatts] to build a prosody tokenizer. The reasons why we adopt the chromagram are as follows:
@@
    Octave-free, thus friendly for unification: There is a significant distribution gap of F0 between speech and singing voice, such as the F0 range of singing voice is much wider [my-svc-paper-workshop]. Chromagram removes octave information between singing voice and speech, making it easier to capture common prosodic patterns across both domains.
@@
    Notation-free, thus friendly for scaling up: Chromagram effectively represents high-level musical and singing voice attributes such as melody and harmony [meinard-book], and can be automatically extracted from audio without manual transcription or music notation, making it suitable for large-scale data processing.
@@
    Robust for diverse audio domains: As a statistical distribution of spectral energy, chromagram is less sensitive to noise and pitch detection errors [yin], resulting in more stable performance. In addition, chromagram provides strong representations for both human voice and non-human sounds, such as instrumental music [meinard-book, musicgen]. These characteristics enhance the flexibility of melody control in downstream music generation tasks, such as enabling instrument-to-sing applications (Section V-F).
The prosody tokenizer consists of five main components (Figure 2(a)): Downsampling Layer (CNN-based), Encoder, Vector Quantization (VQ), Decoder, and Upsampling Layer (CNN-based). Formally, given the codebook $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}]$ with vocabulary size $K$, and taking chromagram features $\bm{x}$ as input, the reconstructed output $\hat{\bm{x}}$ is obtained after passing through these five modules:
@@
We follow [vqgan-cv, rvqgan, maskgct] and adopt factorized (i.e., a low-dimensional code-lookup space) and L2-normalized strategies for VQ learning. We extract 50 Hz chromagram features from raw audio and apply an 8x downsampling, resulting in a prosody tokenizer operating at 6.25 Hz (56.25 bps).
Content-Style Tokenizer We follow the main ideas from recent works in controllable speech generation [vevo, cosyvoice], and aim to develop a content-style tokenizer that can encode linguistic content, prosody, and style, while also achieving timbre disentanglement. Specially, to unify both speech and singing voice, beyond encoding linguistic content and style, as in typical speech generation tasks [vevo, cosyvoice, seedtts, maskgct], we also require the tokenizer to capture melody and fine-grained style information (e.g., F0 and energy details) in singing voice. Motivated by this, when designing our VQ-VAE tokenizer, we not only use hidden features from automatic speech recognition (ASR) models—specifically, from Whisper [whisper]—as reconstruction objectives, but also incorporate chromagram features as an additional reconstruction objective. This compensates for the limitations of Whisper features in modeling melody in singing voice [amphion-svc]. Furthermore, to facilitate the downstream content-style modeling stage, we design the content-style tokenizer to operate at a relatively low frame rate (specifically, 12.5 Hz in our study). This reduces the sequence length for the AR model and thus alleviates its learning burden.
The architecture of the content-style tokenizer is similar to the proposed prosody tokenizer, which is displayed in Figure 2(b). The only difference lies in the input and reconstruction objectives incorporate both chromagram features and Whisper features. We extract 50 Hz chromagram features and 50 Hz Whisper features (which is the encoder output of the pre-trained Whisper model [whisper]) and apply a 4x downsampling, resulting in a content-style tokenizer operating at 12.5 Hz (175 bps).

### v3

@@ I Introduction
Controllable generation of human voice is an important research topic in the field of audio generation [65, 74]. In recent years, remarkable progress has been achieved in speech generation, particularly in zero-shot text-to-speech (TTS) [36, 1, 71, 21], largely propelled by the availability of large-scale training corpora. However, for more challenging domains such as singing voice—which can be regarded as a highly expressive form of speech with strictly constrained prosody (e.g., melody) [^1] This paper adopts a definition in which prosody refers to the suprasegmental attributes of human voice, including pitch, duration, and loudness. We regard melody—which primarily involves pitch and duration (with rhythm viewed as a temporal organization of duration)—as a specific and more constrained form of prosody. [64]—achieving both high generation quality and controllability remains challenging [4, 83].
@@
This paper focuses on unifying controllable speech and singing voice generation. Our motivation is based on the hypothesis that learning for speech and singing voice can be mutually beneficial through unified modeling. Specifically, (1) Speech $\rightarrow$ Singing Voice: the abundance of speech data can supplement the typically smaller singing voice datasets [30, 99], potentially improving the overall quality of generated singing voice [83]. (2) Singing Voice $\rightarrow$ Speech: the inherent expressiveness of singing voice, when incorporated into a unified learning framework, can enhance expressive speech generation [28]. Besides, since singing voices inherently contain structured melodic patterns, learning to model them helps the system enhance its ability to follow melody, thereby improving overall prosody-following capability. This capability is particularly beneficial for tasks such as prosody-preserved speech editing (Section V-B).
@@
Several recent studies have attempted to unify speech and singing voice generation. For instance, UniSyn [41] integrates TTS and Singing Voice Synthesis (SVS) through a task identifier that switches between generation modes, while UniAudio [79] extends this concept to a broader range of audio generation tasks via task-conditioned sequence modeling. Although these approaches represent valuable progress, they still face two fundamental challenges. First, the existing corpora for singing voice generation rely heavily on expert annotations [66, 44, 98, 101], such as detailed music notation with precise alignment between lyrics and musical notes [89, 16, 99]. These datasets are scarce, and their unique data organization is inherently inconsistent with large-scale speech corpora, thereby limiting scalability and unified training. Second, achieving versatile controllability within such a unified system remains nontrivial—such as the independent manipulation of text (or lyrics), prosody (or melody), style (e.g., accent, emotion, or singing techniques [73, 99]), and timbre (i.e., speaker identity) within a single framework.
@@
Voice Conversion (style-converted) source wav, reference wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Singing Voice Conversion (style-converted) $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
@@
Voice Conversion (style-preserved) source wav, reference wav - $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$]
@@
To address these challenges, this paper introduces Vevo2, a versatile and controllable voice generation framework designed to unify speech and singing voice. Vevo2 adopts a two-stage architecture, similar to classic zero-shot speech generation systems [1, 21, 95] (Figure 1): (1) Auto-Regressive (AR) Content-Style Modeling Stage: given a target text and a prosodic source as input, we utilize an AR transformer to generate the content-style tokens, which is prompted by a style reference. (2) Flow-Matching (FM) Acoustic Modeling Stage: given content-style tokens as input, we employ an FM transformer to produce acoustic representations (specifically, Mel spectrograms in our study), which is prompted by a timbre reference. To achieve the desired unification and controllability, Vevo2 introduces the following three key method designs:
1. Unified Speech and Singing Voice Tokenizers. We propose two distinct audio tokenizers: a unified prosody tokenizer and a unified content-style tokenizer. (1) Prosody Tokenizer: A single-codebook VQ-VAE tokenizer [67, 39] operating at 6.25 Hz (56.25 bps), trained to reconstruct the chromagram [51] of raw audio (Figure 2(a)). This tokenizer unifies the modeling of both melody in singing voices and prosody in speech, and can directly extract tokens from the chromagram of raw audio, eliminating the need for expert-annotated music notation and thus greatly improving scalability. Moreover, due to the chromagram’s strong generalization ability across various types of audio data [51, 14], this tokenizer can even capture the melody of non-human sounds such as instrumental music (Section V-F), thereby enabling unique music generation applications, such as humming-to-singing and instrument-to-singing. (2) Content-Style Tokenizer: It employs a similar architecture to the prosody tokenizer, but operates at 12.5 Hz (175 bps) and is trained to reconstruct both chromagram features and Whisper hidden features [58] (Figure 2(b)). It effectively captures linguistic content, melody, and style in both speech and singing voice, while also achieving robust timbre disentanglement (Table III).
@@
3. Multi-objective Alignment. To further improve the text- and prosody-following abilities of the pre-trained model, we introduce a multi-objective post-training task. We propose a composite reward that incorporates both intelligibility and prosody similarity, and adopt the Group Relative Policy Optimization (GRPO) algorithm [60] for training. Experimental results demonstrate that this post-training significantly enhances the model’s controllability with respect to text and melody, and also improves its generalization to out-of-pretraining-distribution data, such as instrumental sounds.
@@
We will release code and model checkpoints of this study at Amphion [^2] https://github.com/open-mmlab/Amphion [94, 42].
@@ II Related Work
Controllable Voice Generation Controllable speech and singing voice generation has consistently remained a significant research focus. (1) In the speech domain, existing works explore control over multiple attributes: linguistic content, which is fundamental to TTS [65, 74] and speech editing tasks [40, 54, 70, 11]; timbre, which is central to VC tasks [63, 57, 36, 2, 95, 46]; and style, which encompasses accent [100, 76, 95], emotion [102, 56, 95], whisper [50], and a broader natural language description [48, 34, 19]. (2) For singing voice, beyond controlling lyrics, timbre, and style—as explored in SVS [44, 96, 98, 97, 29], SVC [33, 52, 59, 91], and singing lyric editing tasks [90, 31, 78]melody control holds particular significance due to the inherent musicality of singing voices [14, 4, 83]. (3) From a unified modeling perspective, previous research has attempted to integrate various speech generation tasks such as TTS and VC [95, 21, 72], or different singing voice generation tasks like SVS and SVC [31, 15]. However, models capable of unifying both speech and singing voice domains remain scarce [79, 41]. Moreover, the potential benefits of joint speech and singing pre-training, along with the versatile control mechanisms for multiple voice attributes within a unified system, remain largely unexplored.
@@
Speech Tokenizer Recent research has emphasized the disentangled speech representation learning, aiming to separately model the factorized speech attributes to reduce learning complexity [35, 36] and achieve more precise controllability [55, 36, 95]. Common speech tokenizers include linguistic content tokenizers [32, 95, 3], content-style tokenizers [6, 1, 71, 21, 95], and prosody tokenizers [35, 36]. However, in pursuing a unified controllable model for both speech and singing generation, existing speech tokenizers exhibit several limitations: As these tokenizers predominantly focus on the speech domain, their effectiveness in handling singing voice data remains unexplored. The design of a unified prosody tokenizer capable of handling both modalities, along with enhancing the melody encoding capabilities of content-style tokenizers [91], remains an important issue for future exploration.
@@
Post-training for Audio Generation Alignment via post-training has demonstrated its effectiveness in the generation of text [53, 5], vision [75, 23], speech [86, 1, 22], music [13], and sound effects [49, 43]. In speech generation, existing works have employed preference alignment to enhance multiple aspects of speech, including intelligibility [1, 22, 93], speaker similarity [1, 22], emotion controllability [1, 24], and overall quality [86, 9, 80]. For developing a unified speech and singing voice generation model, the importance of specific capability elicitation through post-training becomes particularly crucial, given the expanded range of capabilities injected into the pre-trained model. However, the design of appropriate alignment objectives for unified speech and singing voice generation remains an open research problem.
@@ III Methodology
The overall architecture of Vevo2 is a two-stage, versatile voice generation framework that unifies both speech and singing voice (Figure 1). During inference, by flexibly combining different input conditions—such as text, prosodic source, style reference, and timbre reference—Vevo2 supports a wide range of voice-related synthesis, conversion, and editing tasks (Section III-D). Specifically, Vevo2 first encodes linguistic, stylistic, and prosodic cues into discrete content-style tokens through an AR transformer, and then reconstructs Mel spectrograms via an FM transformer. These spectrograms are subsequently converted into waveforms using a Vocos-based vocoder [62]. To achieve the unified speech-singing voice generation and versatile controllability, Vevo2 is built upon several key methodological designs, which are detailed in the following sections.
@@ III-A Unified Speech and Singing Voice Tokenizers
Prosody Tokenizer In order to model both the prosody of speech and the melody of singing voice using a unified method, we choose chromagram feature [51] and design a VQ-VAE tokenizer [67, 39] based on it. Some previous studies of speech generation tried to use fundamental frequency (F0) [55, 36] or the low-frequency band of the mel spectrogram [35] to build a prosody tokenizer. The reasons why we adopt the chromagram are as follows:
@@
    Octave-free, thus friendly for unification: There is a significant distribution gap of F0 between speech and singing voice, such as the F0 range of singing voice is much wider [92]. Chromagram removes octave information between singing voice and speech, making it easier to capture common prosodic patterns across both domains.
@@
    Notation-free, thus friendly for scaling up: Chromagram effectively represents high-level musical and singing voice attributes such as melody and harmony [51], and can be automatically extracted from audio without manual transcription or music notation, making it suitable for large-scale data processing.
@@
    Robust for diverse audio domains: As a statistical distribution of spectral energy, chromagram is less sensitive to noise and pitch detection errors [17], resulting in more stable performance. In addition, chromagram provides strong representations for both human voice and non-human sounds, such as instrumental music [51, 14]. These characteristics enhance the flexibility of melody control in downstream music generation tasks, such as enabling instrument-to-sing applications (Section V-F).
Although a continuous chromagram can be directly used as a conditioning signal, we adopt VQ quantization to obtain compact discrete prosody tokens that are better suited for our sequence modeling pipeline and more robust to acoustic variations (see Appendix for an empirical analysis). The prosody tokenizer consists of five main components (Figure 2(a)): Downsampling Layer (CNN-based), Encoder, Vector Quantization (VQ), Decoder, and Upsampling Layer (CNN-based). Formally, given the codebook $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}]$ with vocabulary size $K$, and taking chromagram features $\bm{x}$ as input, the reconstructed output $\hat{\bm{x}}$ is obtained after passing through these five modules:
@@
We follow [81, 39, 71] and adopt factorized (i.e., a low-dimensional code-lookup space) and L2-normalized strategies for VQ learning. We extract 50 Hz chromagram features from raw audio and apply an 8x downsampling, resulting in a prosody tokenizer operating at 6.25 Hz (56.25 bps).
Content-Style Tokenizer We follow the main ideas from recent works in controllable speech generation [95, 21], and aim to develop a content-style tokenizer that can encode linguistic content, prosody, and style, while also achieving timbre disentanglement. Specially, to unify both speech and singing voice, beyond encoding linguistic content and style, as in typical speech generation tasks [95, 21, 1, 71, 84], we also require the tokenizer to capture melody and fine-grained style information (e.g., F0 and energy details) in singing voice. Motivated by this, when designing our VQ-VAE tokenizer, we not only use hidden features from automatic speech recognition (ASR) models—specifically, from Whisper [58]—as reconstruction objectives, but also incorporate chromagram features as an additional reconstruction objective. This compensates for the limitations of Whisper features in modeling melody in singing voice [91]. Furthermore, to facilitate the downstream content-style modeling stage, we design the content-style tokenizer to operate at a relatively low frame rate (specifically, 12.5 Hz in our study). This reduces the sequence length for the AR model and thus alleviates its learning burden.
The architecture of the content-style tokenizer is similar to the proposed prosody tokenizer, which is displayed in Figure 2(b). The only difference lies in the input and reconstruction objectives incorporate both chromagram features and Whisper features. We extract 50 Hz chromagram features and 50 Hz Whisper features (which is the encoder output of the pre-trained Whisper model [58]) and apply a 4x downsampling, resulting in a content-style tokenizer operating at 12.5 Hz (175 bps).